OpenStudy (anonymous):

Will fan and medal:) The radius, r, of a cylinder with a volume of 28.26y cubic units is found by the equation below. r= (28.26y/3.14y)^1/2 Which statement best describes 28.26y? The volume of the cylinder is 28.26 times the height. The volume of the cylinder is 9 times the height. The volume of the cylinder is 28.26 times the radius. The volume of the cylinder is 9 times the radius.

OpenStudy (dumbcow):

y = height

OpenStudy (jhannybean):

Volume of a cylinder: \(V = \pi r^2 h\)Solving for r we get \[r^2 = \frac{V}{\pi h}\]\[r=\sqrt{\frac{V}{\pi h}}\]So if \(V=28.26y\)... can you figure it out?